A bank offers a savings bond with an interest rate of 8% per year compounded annually. How much money would an individual invest so that the value of the investment after 15 years is $7000?
Answer by checkley71(8403)

7000=X(1+.08)^15
7000=X(1.08)^15
7000=X*3.172
X=7000/3.172
X=$2206.81 IS THE AMOUNT OF THE ORIGINAL INVESTMENT.